OpenStudy (frank):

HARD time understanding why an equation is not seperable. As an example, if the integral of Y dx = xy + c, then why does the derivative of xy = y + x dy/dx ? Also... this textbook says that dy/dx=y + sinx is not seperable. If I take the integral of both sides with respect to dx I get : integral ( dy/dx * dx) = integral (y + sinx) dx which becomes y=yx -cos(x). Im pretty sure my math for the integration is right, yet like the integral of y = yx, it just doesnt reintegrate correctly. Reintegrating becomes : dy/dx=y+xdy/dx+sin(x) What am I missing ? Its driving me crazyyyy

OpenStudy (anonymous):

You ca can't integrate respect to x bouth sides of dy/dx=y + sinx, because y is actualy y(x), and you , at this point at least, have no idea how this function looks like.

OpenStudy (amistre64):

separable means that you can separate it into F(x) dx = G(y) dy

OpenStudy (amistre64):

Y dx = xy + c ^^ you cant factor this into (x) and (y) parts that readily

OpenStudy (frank):

ok I am willing to accept the fact that... You cant integrate it because you dont know what y really is, it could be x^100e^y or whatever crazy. but then how come wolfram alpha says : integral of y dx = yx, but the derivative of yx is not y?

OpenStudy (amistre64):

you need to notate that a little better ...

OpenStudy (frank):

amistre : i know the RULE is both have to be seperated... the thing is I want to know why. I dont want to just know the rule.

OpenStudy (amistre64):

becasue in order to separate it up and apply a separation method ... you have to be able to separate it.

OpenStudy (anonymous):

\(\int ydx=yx+\phi(y)\), that's how it should be

OpenStudy (frank):

wow. so anytime you integrate y with respect to x... that means its not + c but its + a function ?

OpenStudy (amistre64):

there might be some blurring between PDEs and ODEs

OpenStudy (anonymous):

to clear the things, jst follow what @amistre64 said befor: separable means that you can separate it into F(x) dx = G(y) dy

OpenStudy (frank):

well. I can move forward that way, but not knowing what im doing wrong here, to me, messes up my whole foundation. I really appretiate both of your help though.

OpenStudy (amistre64):

\[\int\limits y(x)~ dx = Y(x) + C\] which is what an ordinary differential equation tends to represent

OpenStudy (anonymous):

yes, just: \(\int ydx=xy+C(y)\)

OpenStudy (anonymous):

this is in case if you treat y as just another variable

OpenStudy (amistre64):

\[\frac{d}{dx}[xy(x)]\to y(x)+x\frac{dy}{dx}\]or \[g(x,y(x),\frac{dy}{dx})=y(x)+x\frac{dy}{dx}\] and thats what you are lookinng at when integrating

OpenStudy (amistre64):

\[g(x,y(x),\frac{dy}{dx})=y(x)+x\frac{dy}{dx}\] \[g(x,y(x),\frac{dy}{dx})-y(x)=x\frac{dy}{dx}\] \[\left[g (x,y(x),\frac{dy}{dx})-y(x)\right]dx~=x~dy\]

OpenStudy (amistre64):

that just doesnt look like a very nice "separation" to me

OpenStudy (frank):

is it because the\[ \int\limits\limits g(x,y(x),\frac{dy}{dx}) dx = \int\limits \frac{d}{dx}xy(x)dx - \int\limits xy(x)\] and that last integral would be with respect to nothing... so thats why it wouldnt work?

OpenStudy (frank):

that - should be an = sign. basically im saying, it doesnt work because the d/dx will cancel with another dx and then youd have an intergral with respect to nothing.

OpenStudy (amistre64):

it would prolly be best to simplify it as: \[\left[g (x,y(x),\frac{dy}{dx})-y(x)\right]dx~=x~dy\] \[since:~g(x,y(x),\frac{dy}{dx})=\] \[\left[y(x)+x\frac{dy}{dx}-y(x)\right]dx~=x~dy\] \[\left[x\frac{dy}{dx}\right]dx~=x~dy\] \[x~dy~=x~dy\] which seems rather trivial: but\[\int x~dy=xy+f(y)\]

OpenStudy (amistre64):

let f(y) = 0 to get back to your inital setup if need be